OpenStudy (anonymous):

Convert the complex number 3-3i into its polar representation. A. 3(cos(60°) + isin(60°)) B. 3√2 (cos(135°) + isin(135° )) C. √3 (cos(225°) + isin(225°)) D. 3√2 (cos(315°) + isin(315°))

ganeshie8 (ganeshie8):

\(\large a+bi = r\left(\cos (\theta) + i\sin(\theta)\right)\)

ganeshie8 (ganeshie8):

\(\large r = \sqrt{a^2+b^2}\) \(\theta = \arctan(b/a)\)